OpenStudy (anonymous):

how to find cot 120 I know it is negative

ganeshie8 (ganeshie8):

split it into known angles

OpenStudy (anonymous):

cos120/sin120 ?

ganeshie8 (ganeshie8):

yes ! cos/sin works too ^

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

looks good !

ganeshie8 (ganeshie8):

i would have done like this -

ganeshie8 (ganeshie8):

since the period of cot is 180, i can add/subtract 180 and the value will not change cot(120) = cot(120-180) = cot(-60) = -cot(60) = -sqrt(3)/3

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

If the period is 360 as for for sin and cos and we subtract or add 180 or 90 by how much will the value change

OpenStudy (anonymous):

is it this logic as 90 is 1/4 th of 360 value would be 1/4 th

ganeshie8 (ganeshie8):

the value will NOT change only if we add/subtract the PERIOD of function.

ganeshie8 (ganeshie8):

adding/subtracting 90/180 etc is a different logic : it comes directly from the angle sum/difference identities

ganeshie8 (ganeshie8):

for example : sin(x+90) = sinx cos90 + cosx sin90 = sinx 0 + cosx 1 = cos x

OpenStudy (anonymous):

ohkk :)

OpenStudy (anonymous):

cot(x+90)=?

ganeshie8 (ganeshie8):

cot(x+90) = cos(x+90) / sin(x+90) = -sinx / cosx = -tan x

OpenStudy (anonymous):

oh yes